OpenStudy (anonymous):

f(x) = |x + 1|g(x) = 3x - 2 Find (f*g)(-2) - 7 - 1 1 7

OpenStudy (anonymous):

i am guessing that this is composition \(f\circ g\) not \(f\times g\) am i right?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

what is \(g(-2)\) ?

OpenStudy (anonymous):

\[f(x)=\left| x+1 \right|\]\[g(x)=3x-2\]\[f(g(x))=\left| (3x-2)+1 \right| \]To find f(g(-2)), we simply plug in x = -2 in our f(g(x)) equation.\[f(g(-2))=\left| (3(-2)-2)+1 \right|=?\]Can you solve it from here? That's one way of doing it. The other way is that since we know we are finding f(g(-2)), we must first find g(-2). So plug in -2 in to g(x) and evaluate. Then plug in that value in to f(x) for x and solve to get f(g(-2)). @Annetta_Martin